# Gravitational reduction of the wave function through the quantum theory of motion

### v1

@@ 7 Conclusions
A Bohmian picture for gravity-induced wave function reduction was provided. Through this approach, it is possible to investigate the problem of gravitational reduction of the wave function with concrete concepts such as force and particle motion. Examining the deviation between nearby trajectories in the ensemble led to the condition of equilibrium between quantum and quantum gravitational forces, which gave us the critical width of the transition between the quantum and classical worlds. We also obtained the relation (28), which classifies the different regimes of motion in terms of the ratio of particle mass to the critical mass. In a gravity-dominant regime, where only the quantum gravitational force is considered, we investigated the equation of motion of the particle and found that the particle performs oscillatory motion within a certain range of its quantum distribution. What is discussed as the collapse time of the wave function in standard quantum mechanics is related here to the period of the particle’s oscillatory motion. We showed that this oscillatory motion is not simple and depends on the amplitude of the motion. In other words, particles with wider wave packets take longer to reach the center of the probability distribution and vice versa. But what determines the width of a wave packet or probability distribution is the mass of the particle. The probability distribution of heavier particles is narrower than that of lighter particles. The behavior of particle motion was illustrated in Figs. (3) and (4). Additionally, the motion of the particle was examined in a mixed regime, where both quantum and gravitational forces are present. It was observed that as the mass decreases, the particle moves away from the distribution center due to the quantum force, and the reduction of the corresponding wave function is never observed. Conversely, as the mass of the particle increases, the gravitational regime prevails, and the particle performs oscillatory motion within a range determined by its mass.
@@
For an object, we obtained the critical width using the concepts of Bohmian quantum mechanics. The results are in agreement with those obtained by Diosi in her studies. We showed that for micro objects, one must consider the asymptotic form of the quantum gravitational force ($\sigma_{0}\gg R$) to achieve the correct result for the critical width of the associated wave packet. The equation of motion of an object (Eq. 58) was investigated numerically, revealing a non-simple oscillatory motion in a gravity-dominant regime. Using equation (58), it is possible to check the motion of the object for different initial conditions and parameter values. While the period of motion can be obtained numerically through the equation of motion, we estimated the reduction time through the uncertainty in the quantum gravitational energy of the object to achieve approximate analytical results. As we mentioned earlier, similar to the standard quantum mechanics, we do not need an ensemble of metrics; rather, we require an ensemble of particle trajectories, with their quantum evolution considered in a fixed background. It seems that this approach can be further studied to obtain more comprehensive results.

### v2

@@ 4 Conclusion
A Bohmian perspective for gravity-induced wave function reduction was provided. Through this approach, it is possible to investigate the problem of gravitational reduction of the wave function with concrete concepts such as force and particle motion. Examining the deviation between nearby trajectories in the ensemble led to the condition of equilibrium between quantum and quantum gravitational forces, which is used to obtain the critical width for the transition between the quantum and classical worlds. We also obtained the relation (28), which classifies the different regimes of motion in terms of the ratio of particle mass to the critical mass. In a gravity-dominant regime, where only the quantum gravitational force is considered, we investigated the equation of motion of the particle and found that the particle performs oscillatory motion within a certain range of its quantum distribution. What is discussed as the collapse time of the wave function in standard quantum mechanics is related here to the period of the particle’s oscillatory motion. We showed that this oscillatory motion is not simple and depends on the amplitude of the motion. In other words, particles with wider wave packets take longer to reach the center of the probability distribution and vice versa. The behavior of particle motion was illustrated in Figs. (2) and (3). Additionally, the motion of the particle was examined in a mixed regime, where both quantum and gravitational forces are present. It was observed that as the mass decreases, the particle moves away from the distribution center due to the quantum force, and the reduction of the corresponding wave function is never observed. Conversely, as the mass of the particle increases, the gravitational regime prevails, and the particle performs oscillatory motion within a range determined by its mass.
@@
For an object, we obtained the critical width using the concepts of Bohmian quantum mechanics. The results are in agreement with those obtained by Diosi in her studies. We showed that for micro objects, one must consider the asymptotic form of the quantum gravitational force ($\sigma_{0}\gg R$) to achieve the correct result for the critical width of the associated wave packet. The equation of motion of an object (Eq. 58) was investigated numerically, revealing a non-simple oscillatory motion in a gravity-dominant regime. Using equation (58), it is possible to check the motion of the object for different initial conditions and parameter values. While the period of motion can be obtained numerically through the equation of motion, we estimated the reduction time through the uncertainty in the quantum gravitational energy of the object to achieve approximate analytical results. This approach provides a clearer understanding of the underlying processes, thanks to the deterministic structure of Bohmian quantum mechanics. It seems that this approach can be further studied to obtain more comprehensive results.